OpenStudy (anonymous):

Math help?? (2/3)^2 (3/4)^-2 (2x^6/y^4)^3 (a/5b)^-2 Ill give best response lol.

OpenStudy (anonymous):

Mathway.com

OpenStudy (anonymous):

\(\sf\Huge(\frac{a}{b})^x=\frac{a^x}{b^x}\)

OpenStudy (anonymous):

@iamrobin What if the fraction already has exponents?

OpenStudy (anonymous):

\(\sf\Huge(\frac{a^c}{b^c})^d=(\frac{a^{c \times d}}{b^{c \times d}})\)

OpenStudy (anonymous):

@iamrobin Thank you!

OpenStudy (anonymous):

Don't mention it :)